Using the domain values {-1, 0, 4}, find the range values for relation y=2x-7.

Answer by ewatrrr(24785)


Hi,
y = 2x-7
range values corresponding to the domain values {-1, 0, 4}
y = 2*-1 - 7 = -9
y = 2*0 - 7 = -7
y = 2*4 - 7 = 1
corresponding range values are {-9,-7,1}